THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 109 — 1 February 2011

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Figure 1: Rubidium is detected as a strong absorption line at 780 nm. The spectra (in red) of three Rb-rich stars discovered in the Magellanic Clouds are shown together with one of the ESO/VLT telescopes used in this study and the Large Magellanic Cloud (in the upper right corner). Image credits: ESO/IAC/Anibal García-Hernández.

Editorial

Dear Colleagues, It is my pleasure to present you the 109^{th} issue of the Magellanic Clouds Newsletter. Eclipsing binaries, star clusters, and supernova remnants feature prominently this time. Lucky are those in search for a postdoctoral position, because there is a fantastic opportunity in an exciting research group in beautiful Tasmania. Do check out the MAGMA image release – here’s looking forward to the full cubes! Front pictures and other contributions remain welcome. If more than one picture is received it will still be included (but not necessarily on the front cover). I am sure you are as delighted as I am by the inspiring contribution from Sidney van den Bergh, who has studied the Magellanic Clouds for so long he can probably tell you their proper motion from the constellations they used to be in! The next issue is planned to be distributed on the 1^{st} of April 2011.

Editorially Yours, Jacco van Loon

A strange mènage á trois

Sidney van den Bergh

Dominion Astrophysical Observatory, Herzberg Institute of Astrophysics, National Research Council of Canada, 5071 West Saanich Road, Victoria, BC, V9E 2E7, Canada. email: sidney.vandenbergh@nrc-cnrc.gc.ca

The Magellanic Clouds may have joined our Milky Way system quite recently. The Large Magellanic Cloud turns out to be a remarkably luminous object that is close to the upper luminosity limit of the class of magellanic irregular galaxies.

We are all Copernicans now. So we expect to be living in a typical galaxy in a normal neighborhood. The first of these expectations is fulfilled. Our Milky Way system is a relatively normal giant spiral of Hubble type Sbc, or perhaps a barred giant of type SBbc. However, the Galactic neighborhood is unusual and quite different from what might have been expected. True, the Local Group in which we live is a small poor cluster of galaxies like many others in nearby regions of the Universe. However, the nearest neighbors to our home galaxy exhibit two remarkable peculiarities. For most galaxies, like that in Andromeda ^{1} , the nearest neighbors are early-type galaxies of types E or S0, whereas the more distant companions are late-type objects of types Sc or Ir. However, the Milky Way’s two closest big companions, the Large Magellanic Cloud (LMC) and the Small Magellanic Cloud (SMC) are irregular galaxies. This anomaly suggests ^{2} that the Magellanic Clouds might not be close satellites of the Galaxy, but objects that formed in the outer reaches of the Local Group that just happen to be passing close to the Milky Way system at the present time. Recent calculations ^{3} suggest that there is a ∼72% probability that the Magellanic Clouds were accreted within the last Gyr, and a ∼50% probability that they were accreted together. The second anomaly among the closest companions to our Galaxy is that the LMC is so extraordinarily luminous for a magellanic irregular galaxy. In nearby regions of the Universe there are only two Ir galaxies (NGC 4214 and NGC 4449) that even come close to rivaling the LMC in luminosity. In other words the Large Magellanic Cloud seems to be close to the upper luminosity limit of M_{B} ≃ −18.5 for irregular galaxies. This is important because there is a fundamental morphological difference between spiral and irregular galaxies: spirals all have nuclei, whereas Magellanic irregulars do not. It should be emphasizes that this upper luminosity limit applies only to magellanic irregulars and not to the peculiar chaotic irregular galaxies that might have been formed during the collisions or mergers of massive ancestral galaxies.

In 1969 Erik Holmberg ^{4} searched for the satellites of nearby galaxies on the prints of the Palomar Sky Survey. Surprisingly he found that bright satellite galaxies, like the Magellanic Clouds, are quite rare. This conclusion has recently been strengthened and confirmed by the work of James & Ivory ^{5} and by Lui et al.^{6}. James & Ivory used narrow-band imaging of 143 luminous spiral galaxies, comparable to the Milky Way, to search for star forming companions. They concluded that luminous star-forming satellite galaxies are quite rare and that our home galaxy is unusual, both for the luminosity, and the proximity of its two brightest satellites. A different approach was used by Liu et al. who employed the enormous data-base provided by the Sloan Digital Sky Survey to search host galaxies, with luminosities within ± 0.2 mag of that of the Milky Way for satellite galaxies with luminosities similar to those of the Magellanic Clouds, that are located within a distance of 150 kpc of of their apparent host galaxy. For 22581 Milky Way-like hosts they find that 81% have no satellites as bright as the Magellanic Clouds, 11% have one such satellite, and only 3.5% host two such galxies. As Edwin Hubble ^{7} said many years ago “The fact that the [G]alactic system is a member of a group is a very fortunate accident.” That the Galaxy should have an irregular companion as luminous as the Large Magellanic Cloud is almost a miracle.

References: 1 Einastro, J. et al. Nature, 252, 111-113 (1974) 2 van den Bergh, S. Astron. J, 132, 1571-1574 (2006) 3 Busha, M. T. et al. arXiv: 1011,2203v2 (2010) 4 Holmberg, E. Ark. Astron. 5, 305-343 (1969) 5 James, P. A., & Ivory, C. F. Mon.Not.R.Astron.Soc, in press arXiv:1109.2875 (2010) 6 Liu, L. et al. arXiv: 1011.2255v2 7 Hubble, E. The Realm of the Nebulae, New Haven-Yale University press, p.125 (1936)

Available from arXiv:1012.3492

Refereed Journal Papers

Red supergiant stars in the Large Magellanic Cloud. I. The Period–Luminosity relation

Ming Yang^{1} and B.W. Jiang^{1}

1Department of Astronomy, Beijing Normal University, Beijing 100875, China

From previous samples of red supergiants (RSGs) by various groups, 191 objects are assembled to compose a large sample of RSG candidates in LMC. For 189 of them, the identity as an RSG is verified by their brightness and color indexes in several near- and mid-infrared bands related to the Two Micron All Sky Survey (2MASS) JHK_{s} bands and the Spitzer/IRAC and Spitzer/MIPS bands. From the visual time-series photometric observations by the ASAS and MACHO projects which cover nearly 8–10 years, the period and amplitude of light variation are analyzed carefully using both the phase dispersion minimization and Period04 methods. According to the properties of light variation, these objects are classified into five categories: (1) 20 objects are saturated in photometry or located in crowded stellar field with poor photometric results, (2) 35 objects with too complex variation to have any certain period, (3) 23 objects with irregular variation, (4) 16 objects with semi-regular variation, and (5) 95 objects with long secondary period (LSP) among which 31 have distinguishable short period and 51 have a long period shorter than 3000 days that can be determined with reasonable accuracy. For the semi-regular variables and the LSP variables with distinguishable short period, the period–luminosity (P–L) relation is analyzed in the visual, near-infrared, and mid-infrared bands. It is found that the P–L relation is tight in the infrared bands such as the 2MASS JHK_{s} bands and the Spitzer/IRAC bands, in particular in the Spitzer/IRAC [3.6] and [4.5] bands; meanwhile, the P–L relation is relatively sparse in the V band which may be caused by the inhomogeneous interstellar extinction. The results are compared with others’ P–L relationships for RSGs and the P–L sequences of red giants in LMC.

Accepted for publication in The Astrophysical Journal Available from arXiv:1011.4998

The VMC survey – I. Strategy and first data

M.-R.L. Cioni^{1,2}, G. Clementini^{3}, L. Girardi^{4}, R. Guandalini^{1}, M. Gullieuszik^{5}, B. Miszalski^{1}, M.-I. Moretti^{6}, V. Ripepi^{7}, S. Rubele^{4}, G. Bagheri^{1}, K. Bekki^{8}, N. Cross^{9}, W.J.G. de Blok^{10}, R. de Grijs^{11}, J.P. Emerson^{12}, C.J. Evans^{13}, B. Gibson^{14}, E. Gonzales-Solares^{15}, M.A.T. Groenewegen^{5}, M. Irwin^{15}, V.D. Ivanov^{16}, J. Lewis^{15}, M. Marconi^{7}, J.-B. Marquette^{17,18}, C. Mastropietro^{19}, B. Moore^{20}, R. Napiwotzki^{1}, T. Naylor^{21}, J.M. Oliveira^{22}, M.

Read^{9}, E. Sutorius^{9}, J. Th. van Loon^{22}, M.I. Wilkinson^{23} and P.R. Wood^{24}

1University of Hertfordshire, Physics Astronomy and Mathematics, Hatfield AL10 9AB, United Kingdom 2University Observatory Munich, Scheinerstraße 1, 81679 München, Germany 3INAF, Osservatorio Astronomico di Bologna, Via Ranzani 1, 40127 Bologna, Italy 4INAF, Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, 35122 Padova, Italy 5Royal Observatory of Belgium, Ringlaan 3, 1180 Ukkel, Belgium 6University of Bologna, Department of Astronomy, Via Ranzani 1, 40127 Bologna, Italy 7INAF, Osservatorio Astronomico di Capodimonte, via Moiariello 16, 80131 Napoli, Italy 8ICRAR, M468, University of Western Australia, 35 Stirling Hwy., Crawley 6009, Western Australia 9University of Edinburgh, Institute for Astronomy, Blackford Hill, Edinburgh EH9 3HJ, United Kingdom 10University of Cape Town, Private Bag X3, Rondebosch 7701, South Africa 11Peking University, Kavli Institute for Astronomy and Astrophysics, Yi He Yuan Lu 5, Hai Dian District, Beijing, China 12Queen Mary University of London, Mile End Road, London E1 4NS, United Kingdom 13UK Astronomy Technology Centre, Blackford Hill, Edinburgh EH9 3HJ, United Kingdom 14Centre for Astrophysics, University of Central Lancshire, Preston PR1 2HE, United Kingdom 15University of Cambridge, Institute of Astronomy, Madingley Rd., Cambridge CB3 0HA, United Kingdom 16European Southern Observatory, Av. Alonso de Cordoba 3107, Casilla 19, Santiago, Chile 17UPMC Univ. Paris 06, UMR7095, Institut d’Astrophysique de Paris, 75014 Paris, France 18CNRS, UMR7095, Institut d’Astrophysique de Paris, 75014 Paris, France 19LERMA, Observatoire de Paris, UPMC, CNRS, 61 Av. de l’Observatoire, 75014 Paris, France 20University of Zurich, Institute for Theoretical Physics, 8057 Zurich, Switzerland 21University of Exeter, School of Physics, Stocker Road, Exeter EX4 4QL, United Kingdom 22University of Keele, School of Physical and Geographical Sciences, Staffordshire ST5 5BG, United Kingdom 23University of Leicester, University Road, Leicester LE1 7RH, United Kingdom 24Mount Stromlo Observatory, RSAA, Cotter Road, Weston Creek, ACT 2611, Australia

The new Visual and Infrared Telescope for Astronomy (VISTA) has started operations. Over its first five years it will be collecting data for six Public Surveys, one of which is the near-infrared YJK_{s} VISTA survey of the Magellanic Clouds system (VMC). This survey comprises the Large Magellanic Cloud (LMC), the Small Magellanic Cloud, the Magellanic Bridge connecting the two galaxies and two fields in the Magellanic Stream. This paper provides an overview of the VMC survey strategy and presents first science results. The main goals of the VMC survey are the determination of the spatially-resolved star-formation history and the three-dimensional structure of the Magellanic system. The VMC survey is therefore designed to reach stars as faint as the oldest main sequence turn-off point and to constrain the mean magnitude of pulsating variables such as RR Lyrae stars and Cepheids. This paper focuses on observations of VMC fields in the LMC obtained between November 2009 and March 2010. These observations correspond to a completeness of 7% of the planned LMC fields. The VMC data are comprised of multi-epoch observations which are executed following specific time constraints. The data were reduced using the VISTA Data Flow System pipeline with source catalogues, including astrometric and photometric corrections, produced and made available via the VISTA Science Archive. The VMC data will be released to the astronomical community following the European Southern Observatory’s Public Survey policy. The analysis of the data shows that the sensitivity in each wave band agrees with expectations. Uncertainties and completeness of the data are also derived. The first science results, aimed at assessing the scientific quality of the VMC data, include an overview of the distribution of stars in colour–magnitude and colour–colour diagrams, the detection of planetary nebulae and stellar clusters, and the K_{s} band light-curves of variable stars. The VMC survey represents a tremendous improvement, in spatial resolution and sensitivity, on previous panoramic observations of the Magellanic system in the near-infrared, providing a powerful complement to deep observations at other wavelengths. Accepted for publication in Astronomy & Astrophysics Available from arXiv:1012.5193 and from http://star.herts.ac.uk/∼vmc/

Dense clumps in giant molecular clouds in the Large Magellanic Cloud: Density and temperature derived from ^{13}CO(J = 3 − 2) observations

Tetsuhiro Minamidani^{1,2}, Takanori Tanaka^{3}, Yoji Mizuno^{3}, Norikazu Mizuno^{4}, Akiko Kawamura^{3}, Toshikazu Onishi^{5}, Tetsuo Hasegawa^{4}, Ken’ichi Tatematsu^{4}, Tatsuya Takekoshi^{2}, Kazuo Sorai^{1,2}, Nayuta Moribe^{3}, Kazufumi Torii^{3}, Takeshi Sakai^{6}, Kazuyuki Muraoka^{5}, Kunihiko Tanaka^{7}, Hajime Ezawa^{4}, Kotaro Kohno^{6}, Sungeun Kim^{8},

Mónica Rubio^{9} and Yasuo Fukui^{3}

1Department of Physics, Faculty of Science, Hokkaido University, N10W8, Kita-ku, Sapporo 060-0810, Japan 2Department of Cosmosciences, Graduate School of Science, Hokkaido University, N10W8, Kita-ku, Sapporo 060-0810, Japan 3Department of Astrophysics, Nagoya University, Furo-cho, Chikusa-ku, Nagoya 464-8602, Japan 4National Astronomical Observatory of Japan, Mitaka, Tokyo 181-8588, Japan 5Department of Physical Science, Osaka Prefecture University, Gakuen 1-1, Sakai, Osaka 599-8531, Japan 6nstitute of Astronomy, The University of Tokyo, 2-21-1 Osawa, Mitaka, Tokyo 181-0015, Japan 7Institute of Science and Technology, Keio University, 4-14-1 Hiyoshi, Yokohama, Kanagawa 223-8522, Japan 8Astronomy & Space Science Department, Sejong University, 98 Kwangjin-gu, Kunja-dong, Seoul, 143-747, Korea 9Departament de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile

In order to precisely determine temperature and density of molecular gas in the Large Magellanic Cloud, we made observations of optically thin ^{13}CO(J = 3 − 2) transition by using the ASTE 10m telescope toward 9 peaks where 12CO(J = 3 − 2) clumps were previously detected with the same telescope. The molecular clumps include those in giant molecular cloud (GMC) Types I (with no signs of massive star formation), II (with H ii regions only), and III (with H ii regions and young star clusters). We detected ^{13}CO(J = 3 − 2) emission toward all the peaks and found that their intensities are 3–12 times lower than those of ^{12}CO(J = 3 − 2). We determined the intensity ratios of 12CO(J = 3 − 2) to ^{13}CO(J = 3 − 2), R_{3−2}^{12/13}, and ^{13}CO(J = 3 − 2) to ^{13}CO(J = 1 − 0), R_{3−2/1−0}^{13}, at 45′′ resolution. These ratios were used for radiative transfer calculations in order to estimate temperature and density of the clumps. The parameters of these clumps range kinetic temperature T_{kin} = 15–200 K, and molecular hydrogen gas density n(H_{2}) = 8 × 10^{2}–7 × 10^{3} cm^{−3}. We confirmed that the higher density clumps show higher kinetic temperature and that the lower density clumps lower kinetic temperature at a better accuracy than in the previous work. The kinetic temperature and density increase generally from a Type I GMC to a Type III GMC. We interpret that this difference reflects an evolutionary trend of star formation in molecular clumps. The R_{3−2/1−0}^{13} and kinetic temperature of the clumps are well correlated with Hα flux, suggesting that the heating of molecular gas n(H_{2}) = 10^{3}–10^{4} cm^{−3} can be explained by stellar FUV photons.

Accepted for publication in The Astronomical Journal Available from arXiv:1012.5037

NGC 1866: a milestone for understanding the chemical evolution of stellar populations in the LMC

A. Mucciarelli ^{1}, S. Cristallo^{2}, E. Brocato^{3}, L. Pasquini^{4}, O. Straniero^{3}, E. Caffau^{5,6}, G. Raimondo^{3}, A. Kaufer^{7}, I. Musella^{8}, V. Ripepi^{8}, M. Romaniello^{4} and A. Walker^{9}

1Astronomy Department, Bologna University, Italy 2Departamento de Física Teórica y del Cosmos, Universidad de Granada, Spain 3INAF – Teramo, Italy 4ESO – Garching, Germany 5GEPI – Paris, France 6ZAH – Landessternwarte Heidelberg, Germany 7ESO – Santiago, Chile 8INAF – Napoli, Italy 9Cerro Tololo Inter-American Observatory, Chile

We present new FLAMES@VLT spectroscopic observations of 30 stars in the field of the LMC stellar cluster NGC 1866. NGC 1866 is one of the few young and massive globular cluster that is close enough so that its stars can be individually studied in detail. Radial velocities have been used to separate stars belonging to the cluster and to the LMC field and the same spectra have been used to derive chemical abundances for a variety of elements, from [Fe/H] to the light (i.e. Na, O, Mg, ...) to the heavy ones. The average iron abundance of NGC 1866 turns out to be [Fe/H] = −0.43±0.01 dex (with a dispersion of 0.04 dex), from the analysis of 14 cluster-member stars. Within our uncertainties, the cluster stars are homogeneous, as far as chemical composition is concerned, independent of the evolutionary status. The observed cluster stars do not show any sign of the light elements ’anti-correlation’ present in all the Galactic globular clusters so far studied, and also found in the old LMC stellar clusters. A similar lack of anti-correlations has been detected in the massive intermediate-age LMC clusters, indicating a different formation/evolution scenario for the LMC massive clusters younger than ∼ 3 Gyr with respect to the old ones. Also opposite to the Galactic globulars, the chemical composition of the older RGB field stars and of the young post-MS cluster stars show robust homogeneity suggesting a quite similar process of chemical evolution. The field and cluster abundances are in agreement with recent chemical analysis of LMC stars, which show a distinctive chemical pattern for this galaxy with respect to the Milky Way. We discuss these findings in light of the theoretical scenario of chemical evolution of the LMC.

Accepted for publication in MNRAS Available from arXiv:1012.1476

The sparsest clusters with O stars

J.B. Lamb^{1}, M.S. Oey^{1}, J.K. Werk^{1,2} and L.D. Ingleby^{1}

1Department of Astronomy, University of Michigan, 500 Church St., Ann Arbor, MI 48109, USA 2Department of Astronomy, Columbia University, 550 West 120^{th} St., New York, NY 10027, USA

There is much debate on how high-mass star formation varies with environment, and whether the sparsest star-forming environments are capable of forming massive stars. To address this issue, we have observed eight apparently isolated OB stars in the SMC using HST’s Advanced Camera for Surveys. Five of these objects appear as isolated stars, two of which are confirmed to be runaways. The remaining three objects are found to exist in sparse clusters, with < 10 companion stars revealed, having masses of 1–4 M_{⊙}. Stochastic effects dominate in these sparse clusters, so we perform Monte Carlo simulations to explore how our observations fit within the framework of empirical, galactic cluster properties. We generate clusters using a simplistic −2 power-law distribution for either the number of stars per cluster (N_{⋆}) or cluster mass (M_{cl}). These clusters are then populated with stars randomly chosen from a Kroupa IMF. We find that simulations with cluster lower-mass limits of M_{cl,lo} > 20 M_{⊙} and N_{⋆,lo} > 40 match best with observations of SMC and Galactic OB star populations. We examine the mass ratio of the second-most massive and most massive stars (m_{max,2}/m_{max}), finding that our observations all exist below the 20^{th} percentile of our simulated clusters. However, all of our observed clusters lie within the parameter space spanned by the simulated clusters, although some are in the lowest 5^{th} percentile frequency. These results suggest that clusters are built stochastically by randomly sampling stars from a universal IMF with a fixed stellar upper-mass limit. In particular, we see no evidence to suggest a m_{max}–M_{cl} relation. Our results may be more consistent with core accretion models of star formation than with competitive accretion models, and they are inconsistent with the proposed steepening of the integrated galaxy IMF (IGIMF).

Published in The Astrophysical Journal Available from arXiv:1010.5273

The outer rings of SN 1987A

Anestis Tziamtzis^{1}, Peter Lundqvist^{1}, Per Gröningsson^{1} and Soroush Nasoudi-Shoar^{2}

1Stockholm Observatory, AlbaNova Science Center, Department of Astronomy, SE-106 91 Stockholm, Sweden 2Argelander-Institut für Astronomie, Auf dem Hügel 71 D-53121 Bonn, Germany

We investigate the physical properties and structure of the outer rings of SN 1987A to understand their formation and evolution. We used low resolution spectroscopy from VLT/FORS1 and high resolution spectra from VLT/UVES to estimate the physical conditions in the outer rings, using nebular analysis for emission lines such as [O ii], [O iii], [N ii], and [S ii]. We also measured the velocity at two positions of the outer rings to test a geometrical model for the rings. Additionally, we used data from the HST science archives to check the evolution of the outer rings of SN 1987A for a period that covers almost 11 years. We measured the flux in four different regions, two for each outer ring. We chose regions away from the two bright neighbouring stars and as far as possible from the inner ring and created light curves for the emission lines of [O iii], Hα, and [N ii]. The light curves display a declining behaviour, which is consistent with the initial supernova flash powering of the outer rings. The electron density of the emitting gas in the outer rings, as estimated by nebular analysis from the [O ii] and [S ii] lines, is < ∼3 × 10^{3} cm^{−3}, has not changed over the last ∼ 15 years, and also the [N ii] temperature remains fairly constant at ∼ 1.2 × 10^{4} K. We find no obvious difference in density and temperature for the two outer rings. The highest density, as estimated from the decay of Hα, could, however, be ∼ 5 × 10^{3} cm^{−3}, and since the decay is somewhat faster in the southern outer ring than it is in the northern, the highest density in the outer rings may be found in the southern outer ring. For an assumed distance of 50 kpc to the supernova, the distance between the supernova and the closest parts of the outer rings could be as short as ∼ 1.7 × 10^{18} cm. Interaction between the supernova ejecta and the outer rings could therefore start in less than ∼ 20 years. We do not expect the outer rings to show the same optical display as the equatorial ring when this happens. Instead soft X-rays should provide a better way of observing the ejecta–outer rings interaction.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1008.3387

Very long-term optical variability of High Mass X-ray Binaries in the SMC

A.F. Rajoelimanana^{1,2}, P.A. Charles^{1,3} and A. Udalski^{4}

1South African Astronomical Observatory, P.O. Box 9, Observatory, 7935, South Africa 2University of Cape Town, Private Bag X3, Rondebosch, 7701, South Africa 3School of Physics and Astronomy, Southampton University, Southampton SO17 1BJ, UK 4Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warsaw, Poland

We have studied the very long-term temporal properties of the optical emission from Be X-ray binaries (BeX) in the Small Magellanic Cloud over a ∼ 16 yr baseline, using light curves from the MACHO and OGLE databases. All the BeX in our sample display superorbital variations, many of them quasi-periodic on timescales of ∼ 200–3000 d. These long-term variations are believed to be related to the formation and depletion of the circumstellar disc around the Be star and we compare and contrast their behaviour with that of the LMC’s prototypical BeX, A 0538−66. The great majority of sources show a correlation of outburst amplitude with brightness (the opposite to that seen in A 0538−66) although the amplitudes are mostly small (< 0.1 mag). We suggest this is an orbital inclination effect. In addition, we have also detected many of their optical orbital periodicities, visible as a series of precisely regular outbursts. Furthermore, the amplitude of these periodic outbursts can vary through the long-term superorbital cycle, and we discuss mechanisms which can produce this effect, as well as examining an apparent correlation between these periodicities. As a by-product of this variation survey we have compiled a list of all the reported SMC BeX orbital and superorbital periodicities at optical and X-ray wavelengths.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1012.4610

A panchromatic view of NGC 602: Time-resolved star formation with the Hubble and Spitzer Space Telescopes

Lynn Redding Carlson^{1}, Marta Sewiło^{2}, Margaret Meixner^{3}, Krista Alexandra Romita^{4}, Barbara Whitney^{5}, J.L. Hora^{6}, M. Cignoni^{7}, E. Sabbi^{3}, A. Nota^{3}, M. Sirianni^{8}, L.J. Smith^{3}, K. Gordon^{3}, B. Babler^{9}, S. Bracker^{9}, J.S.

Gallagher, III^{9}, M. Meade^{9}, K. Misselt^{10}, A. Pasquali^{11} and B. Shiao^{3}

1Leiden Observatory, The Netherlands 2Johns Hopkins University, USA 3Space Telescope Science Institute, USA 4University of Florida, USA 5Space Science Institute, USA 6Harvard/CfA, USA 7Università degli Studi di Bologna, Italy 8ESA/ESTEC, The Netherlands 9University of Wisconsin, USA 10Steward Observatory / University of Arizona, USA 11MPIA, Germany

We present the photometric catalogs for the star-forming cluster NGC 602 in the wing of the Small Magellanic Cloud covering a range of wavelengths from optical (HST/ACS F555W, F814W and SMARTS/ANDICAM V, I) to infrared Spitzer/IRAC 3.6, 4.5, 5.8, and 8 μm and MIPS 24 μm). Combining this with IRSF (InfraRed Survey Facility) nearinfrared photometry (J, H, K_{s}), we compare the young main sequence (MS) and pre-main sequence (PMS) populations prominent in the optical with the current young stellar object (YSO) populations revealed by the infrared (IR). We analyze the MS and PMS population with isochrones in color–magnitude diagrams to derive ages and masses. The optical data reveal ∼ 565 PMS candidates, low mass Stage iii YSOs. We characterize ∼ 40 YSOs by fitting their spectral energy distributions (SEDs) to a grid of models (Robitaille et al. 2007) to derive luminosities, masses and evolutionary phase (Stage i–iii). The higher resolution HST images reveal that ∼ 70% of the YSO candidates are either multiples or protoclusters. For YSOs and PMS sources found in common, we find a consistency in the masses derived. We use the YSO mass function to derive a present-day star-formation rate of ∼ 0.2–1.0 M_{⊙} yr^{−1} kpc^{−2}, similar to the rate derived from the optical star formation history suggesting a constant star formation rate for this region. We demonstrate a progression of star formation from the optical star cluster center to the edge of the star forming dust cloud. We derive lifetimes of a few 10^{5} years for the YSO Stages i and ii.

Accepted for publication in ApJ Available from arXiv:1012.3406

The OGLE view of microlensing towards the Magellanic Clouds. III. Ruling out sub-solar MACHOs with the OGLE-III LMC data

Ł. Wyrzykowski^{1,3}, S. Kozłowski^{2,3}, J. Skowron^{2,3}, A. Udalski^{3}, M.K. Szymański^{3}, M. Kubiak^{3}, G. Pietrzyński^{3,4}, I. Soszyński^{3}, O. Szewczyk^{3,4}, K. Ulaczyk^{3} and R. Poleski^{3}

1Institute of Astronomy, University of Cambridge, UK 2Department of Astronomy, The Ohio State University, USA 3Warsaw University Astronomical Observatory, Poland 4Universidad de Concepción, Departamento de Física, Astronomy Group, Chile

In the third part of the series presenting the Optical Gravitational Lensing Experiment (OGLE) microlensing studies of the dark matter halo compact objects (MACHOs) we describe results of the OGLE-III monitoring of the Large Magellanic Cloud (LMC). This unprecedented data set contains almost continuous photometric coverage over 8 years of about 35 million objects spread over 40 square degrees. We report a detection of two candidate microlensing events found with the automated pipeline and an additional two, less probable, candidate events found manually. The optical depth derived for the two main candidates was calculated following a detailed blending examination and detection efficiency determination and was found to be τ = (0.16 ± 0.12) × 10^{−7}.

If the microlensing signal we observe originates from MACHOs it means their masses are around 0.2 M_{⊙} and they compose only f = 3 ± 2 per cent of the mass of the Galactic Halo. However, the more likely explanation of our detections does not involve dark matter compact objects at all and rely on natural effect of self-lensing of LMC stars by LMC lenses. In such a scenario we can almost completely rule out MACHOs in the sub-solar mass range with an upper limit at f < 7 per cent reaching its minimum of f < 4 per cent at M = 0.1 M_{⊙}. For masses around M = 10 M_{⊙} the constraints on the MACHOs are more lenient with f ∼ 20 per cent. Owing to limitations of the survey there is no reasonable limit found for heavier masses, leaving only a tiny window of mass spectrum still available for dark matter compact objects.

Accepted for publication in MNRAS Available from arXiv:1012.1154 and from http://www.ast.cam.ac.uk/∼wyrzykow/DATA LMC O3/

First VLBI detection of the radio remnant of supernova 1987A: Evidence for small-scale features

C.-Y. Ng^{1,2}, T.M. Potter^{3}, L. Staveley-Smith^{3}, S. Tingay^{4}, B.M. Gaensler^{2}, C. Phillips^{5}, A.K. Tzioumis^{5} and G. Zanardo^{3}

1McGill University, Canada 2The University of Sydney, Australia 3The University of Western Australia, Australia 4Curtin University, Australia 5ATNF, Australia

We present a detailed analysis of the first very long baseline interferometry (VLBI) detection of the radio remnant of supernova 1987A. The VLBI data taken in 2007 and 2008 at 1.4 and 1.7 GHz, respectively, provide images sensitive to angular scales from 0.1′′ to 0.7′′, the highest resolution to date at radio frequencies. The results reveal two extended lobes with an overall morphology consistent with observations at lower resolutions. We find evidence of small-scale features in the radio shell, which possibly consist of compact clumps near the inner surface of the shell. These features have angular extent smaller than 0.2′′ and contribute less than 13% of the total remnant flux density. No central source is detected in the VLBI images. We place a 3-σ flux density limit of 0.3 mJy on any pulsar or pulsar wind nebula at 1.7 GHz.

Accepted for publication in ApJL Available from arXiv:1012.5092

The distance to the massive eclipsing binary LMC-SC1-105 in the Large Magellanic Cloud

A.Z. Bonanos^{1}, N. Castro^{1}, L.M. Macri^{2} and R.-P. Kudritzki^{3,4}

1Institute of Astronomy & Astrophysics, National Observatory of Athens, I. Metaxa & Vas. Pavlou St., P. Penteli, 15236 Athens, Greece 2George P. and Cynthia Woods Mitchell Institute for Fundamental Physics and Astronomy, Department of Physics & Astronomy, Texas A&M University, 4242 TAMU, College Station, TX 77843-4242, USA 3Institute for Astronomy, University of Hawai’i, 2680 Woodlawn Dr., Honolulu, Hawai’i 96822, USA 4Max-Planck-Institute for Astrophysics, Karl-Schwarzschild Str. 1, D-85741 Garching, Germany

This paper presents the first distance measurement to the massive, semi-detached, eclipsing binary LMC-SC1-105, located in the LH 81 association of the Large Magellanic Cloud (LMC). Previously determined parameters of the system (by Bonanos 2009) are combined with new near-infrared photometry and a new temperature analysis to constrain the reddening toward the system, and determine a distance of 50.6 ± 1.6 kpc (corresponding to a distance modulus of 18.5 ± 0.07 mag), in agreement with previous eclipsing binary measurements. This is the sixth distance measurement to an eclipsing binary in the LMC, although the first to an O-type system. We thus demonstrate the suitability of O-type eclipsing binaries as distance indicators. We suggest using bright, early-type eclipsing binaries to measure distances along different sight lines, as an independent way to map the depth of the LMC and resolve the controversy about its 3-dimensional structure.

Submitted to ApJL Available from arXiv:1101.1953

Can rotation explain the multiple main sequence turn-offs of Magellanic Cloud star clusters?

Léo Girardi^{1}, Patrick Eggenberger^{2} and Andrea Miglio^{3}

1Osservatorio Astronomico di Padova – INAF, Italy 2Observatoire de Genève, Université de Genève, Switzerland 3Institut d’Astrophysique et de Géophysique de l’Université de Liège, Belgium

Many intermediate age star clusters in the Magellanic Clouds present multiple main sequence turn-offs (MMSTO), which challenge the classical idea that star formation in such objects took place over short timescales. It has been recently suggested that the presence of fast rotators among main sequence stars could be the cause of such features (Bastian & de Mink 2009), hence relaxing the need for extended periods of star formation. In this letter, we compute evolutionary tracks and isochrones of models with and without rotation. We find that, for the same age and input physics, both kinds of models present turn-offs with an almost identical position in the colour–magnitude diagrams. As a consequence, a dispersion of rotational velocities in coeval ensembles of stars could not explain the presence of MMSTOs. We construct several synthetic colour–magnitude diagrams for the different kinds of tracks and combinations of them. The models that best reproduce the morphology of observed MMSTOs are clearly those assuming a significant spread in the stellar ages – as long as ∼ 400 Myr – added to a moderate amount of convective core overshooting. Only these models produce the detailed “golf club” shape of observed MMSTOs. A spread in rotational velocities alone cannot do anything similar. We also discuss models involving a mixture of stars with and without overshooting, as an additional scenario to producing MMSTOs with coeval populations. We find that they produce turn-offs with a varying extension in the CMD direction perpendicular to the lower main sequence, which are clearly not present in observed MMSTOs.

Accepted for publication in MNRAS Letters Available from arXiv:1101.1880

A classical Cepheid in a LMC eclipsing binary: evidence of shortcomings in current stellar evolutionary models?

Santi Cassisi^{1} and Maurizio Salaris^{2}

1INAF – Astronomical Observatory of Teramo, Italy 2ARI, Liverpool John Moores University, UK

The recent discovery and analysis of a classical Cepheid in the well detached, double-lined, eclipsing binary OGLE-LMC-CEP0227, has provided the first determination of the dynamical mass of a classical Cepheid variable to an unprecedented 1% accuracy. We show here that modern stellar evolution models widely employed to study Galactic and extragalactic stellar systems, are able to match simultaneously mass and radius (and effective temperature) of the two components with a single value for the age of the system, without any specific fine-tuning, assuming the typical metallicity of LMC Cepheids. Our conclusion is that there is no obvious discrepancy between dynamical and evolutionary masses for the Cepheid star in this system, contrary to previous claims of an overestimate of the Cepheid mass by stellar evolution theory.

Accepted for publication in ApJ Letters

Observational constraints on superbubble X-ray energy budgets

A.E. Jaskot^{1}, D.K. Strickland^{2}, M.S. Oey^{1}, Y.-H. Chu^{3} and G. García-Segura^{4}

1University of Michigan, Dept. of Astronomy, 830 Dennison Bldg., Ann Arbor, MI 48109, USA 2Department of Physics and Astronomy, The Johns Hopkins University, 3400 N. Charles St., Baltimore, MD 21218, USA 3University of Illinois, Dept. of Astronomy, Urbana, IL 61801, USA 4Instituto de Astronomía-UNAM, Apartado Postal 877, Ensenada, 22800 Baja California, México

The hot, X-ray-emitting gas in superbubbles imparts energy and enriched material to the interstellar medium (ISM) and generates the hot ionized medium, the ISM’s high-temperature component. The evolution of superbubble energy budgets is not well understood, however, and the processes responsible for enhanced X-ray emission in superbubbles remain a matter of debate. We present Chandra ACIS-S observations of two X-ray-bright superbubbles in the Large Magellanic Cloud (LMC), DEM L50 (N 186) and DEM L152 (N 44), with an emphasis on disentangling the true superbubble X-ray emission from non-related diffuse emission and determining the spatial origin and spectral variation of the X-ray emission. An examination of the superbubble energy budgets shows that on the order of 50% of the X-ray emission comes from regions associated with supernova remnant (SNR) impacts. We find some evidence of mass-loading due to swept-up clouds and metallicity enrichment, but neither mechanism provides a significant contribution to the X-ray luminosities. We also find that one of the superbubbles, DEM L50, is likely not in collisional ionization equilibrium. We compare our observations to the predictions of the standard Weaver et al. model and to 1-D hydrodynamic simulations including cavity supernova impacts on the shell walls. Our observations show that mass-loading due to thermal evaporation from the shell walls and SNR impacts are the dominant source of enhanced X-ray luminosities in superbubbles. These two processes should affect most superbubbles, and their contribution to the X-ray luminosity must be considered when determining the energy available for transport to the ISM.

Accepted for publication in ApJ Available from arXiv:1101.0280

Refined neutron-star mass determinations for six eclipsing X-ray pulsar binaries

Meredith L. Rawls^{1}, Jerome A. Orosz^{2}, Jeffrey E. McClintock^{3}, Manuel A. P. Torres^{3}, Charles D. Bailyn^{4} and Michelle M. Buxton^{4}

1Department of Astronomy, New Mexico State University, USA 2Department of Astronomy, San Diego State University, USA 3Harvard-Smithsonian Center for Astrophysics, USA 4Department of Astronomy, Yale University, USA

We present an improved method for determining the mass of neutron stars in eclipsing X-ray pulsar binaries and apply the method to six systems, namely Vela X-1, 4U 1538−52, SMC X-1, LMC X-4, Cen X-3, and Her X-1. In previous studies to determine neutron star mass, the X-ray eclipse duration has been approximated analytically by assuming the companion star is spherical with an effective Roche lobe radius. We use a numerical code based on Roche geometry with various optimizers to analyze the published data for these systems, which we supplement with new spectroscopic and photometric data for 4U 1538−52. This allows us to model the eclipse duration more accurately and thus calculate an improved value for the neutron star mass. The derived neutron star mass also depends on the assumed Roche lobe filling factor beta of the companion star, where β = 1 indicates a completely filled Roche lobe. In previous work a range of β between 0.9 and 1.0 was usually adopted. We use optical ellipsoidal lightcurve data to constrain beta. We find neutron star masses of 1.77 ± 0.08 M_{⊙} for Vela X-1, 0.87 ± 0.07 M_{⊙} for 4U 1538−52 (eccentric orbit), 1.00 ± 0.10 M_{⊙} for 4U 1538−52 (circular orbit), 1.04 ± 0.09 M_{⊙} for SMC X-1, 1.29 ± 0.05 M_{⊙} for LMC X-4, 1.49 ± 0.08 M_{⊙} for Cen X-3, and 1.07 ± 0.36 M_{⊙} for Her X-1. We discuss the limits of the approximations that were used to derive the earlier mass determinations, and we comment on the implications our new masses have for observationally refining the upper and lower bounds of the neutron star mass distribution.

Accepted for publication in The Astrophysical Journal Available from arXiv:1101.2465

Dusty blastwaves of two young LMC supernova remnants: Constraints on postshock compression

Brian J. Williams^{1}, Kazimierz J. Borkowski^{1}, Stephen P. Reynolds^{1}, Parviz Ghavamian^{2}, John C. Raymond^{3}, Knox

S. Long^{2}, William P. Blair^{4}, Ravi Sankrit^{5}, R. Chris Smith^{6}, Sean Points^{6}, P. Frank Winkler^{7} and Sean P. Hendrick^{8}

1North Carolina State University, USA 2STScI, USA 3Harvard-Smithsonian Center for Astrophysics, USA 4Johns Hopkins University, USA 5SOFIA/USRA, USA 6CTIO, Chile 7Middlebury College, USA 8Millersville University, USA

We present results from mid-IR spectroscopic observations of two young supernova remnants (SNRs) in the Large Magellanic Cloud (LMC) done with the Spitzer Space Telescope. We imaged SNRs B 0509−67.5 and B 0519−69.0 with Spitzer in 2005, and follow-up spectroscopy presented here confirms the presence of warm, shock heated dust, with no lines present in the spectrum. We use model fits to Spitzer IRS data to estimate the density of the postshock gas. Both remnants show asymmetries in the infrared images, and we interpret bright spots as places where the forward shock is running into material that is several times denser than elsewhere. The densities we infer for these objects depend on the grain composition assumed, and we explore the effects of differing grain porosity on the model fits. We also analyze archival XMM-Newton RGS spectroscopic data, where both SNRs show strong lines of both Fe and Si, coming from ejecta, as well as strong O lines, which may come from ejecta or shocked ambient medium. We use model fits to IRS spectra to predict X-ray O line strengths for various grain models and values of the shock compression ratio. For 0509−67.5, we find that compact (solid) grain models require nearly all O lines in X-ray spectra to originate in reverse-shocked ejecta. Porous dust grains would lower the strength of ejecta lines relative to those arising in the shocked ambient medium. In 0519−69.0, we find significant evidence for a higher than standard compression ratio of 12, implying efficient cosmic-ray acceleration by the blast wave. A compact grain model is favored over porous grain models. We find that the dust-to-gas mass ratio of the ambient medium is significantly lower than what is expected in the ISM.

Accepted for publication in The Astrophysical Journal Available from arXiv:1101.1250

Constraining the orbital history of the Magellanic Clouds: A new bound scenario suggested by the tidal origin of the Magellanic Stream

Jonathan Diaz^{1} and Kenji Bekki^{1}

1International Centre for Radio Astronomy Research (ICRAR) – University of Western Australia, Australia

Bound orbits have traditionally been assigned to the Large and Small Magellanic Clouds (LMC and SMC, respectively) in order to provide a formation scenario for the Magellanic Stream (MS) and its Leading Arm (LA), two prominent neutral hydrogen (H i) features connected to the LMC and SMC. However, Hubble Space Telescope (HST) measurements of the proper motions of the LMC and SMC have challenged the plausibility of bound orbits, causing the origin of the MS to re-emerge as a contested issue. We present a new tidal model in which structures resembling the bifurcated MS and elongated LA are able to form in a bound orbit consistent with the HST proper motions. The LMC and SMC have remained bound to each other only recently in our model despite being separately bound to the Milky Way for more than 5 Gyr. We find that the MS and LA are able to form as a consequence of LMC-dominated tidal stripping during the recent dynamical coupling of the LMC and SMC. Our orbital model depends on our assumption that the Milky Way has a constant circular velocity of v_{cir} = 250 km s^{−1} up to 160 kpc, which implies a massive isothermal halo that is not completely rejected by observations.

Accepted for publication in MNRAS Available from arXiv:1101.2500

Influence of the environment on polycyclic aromatic hydrocarbon emission in star-forming regions

V. Lebouteiller^{1,2}, J. Bernard-Salas^{1,3}, D.G. Whelan^{4}, B. Brandl^{5}, F. Galliano^{2}, V. Charmandaris^{6,7,8}, S. Madden^{2} and D. Kunth^{9}

1Center for Radiophysics and Space Research, Cornell University, Space Sciences Building, Ithaca, NY 14853-6801, USA 2Laboratoire AIM, CEA/DSM-CNRS-Université Paris Diderot DAPNIA/Service d’Astrophysique Bât. 709, CEA-Saclay, F-91191 Gif-sur-Yvette Cédex, France 3IAS, Bât. 121, Université Paris-Sud, 91435 Orsay, France 4Department of Astronomy, University of Virginia, P.O. Box 400325, Charlottesville, VA 22904-4325, USA 5Leiden Observatory, Leiden University, P.O. Box 9513, 2300 RA Leiden, The Netherlands 6Department of Physics and Institute of Theoretical & Computational Physics, University of Crete, GR-71003 Heraklion, Greece 7IESL/Foundation for Research and Technology-Hellas, GR-71110 Heraklion, Greece 8Chercheur Associé, Observatoire de Paris, F-75014 Paris, France 9Institut d’Astrophysique de Paris (IAP), 98bis boulevard Arago, 75014 Paris, France

We investigate the emission properties of polycyclic aromatic hydrocarbons (PAHs) in various metallicity environments with the Infrared Spectrograph on board Spitzer. Local giant H ii regions are used as references as they enable access to the distinct interstellar medium components that contribute to the mid-infrared spectrum of star-forming galaxies: photodissociation regions (PDRs), photoionized gas, stellar clusters, and embedded regions. Three objects are considered: NGC 3603 in the Milky Way, 30 Doradus in the Large Magellanic Cloud, and N 66 in the Small Magellanic Cloud. From the variations of the PAH/14 μm ratio, we find that PAHs are destroyed in the ionized gas for a radiation field such that [Ne iii]/[Ne ii] > 3. From the variations of the PAH/Hα ratio, we find that the PAH emission sources in the giant H ii regions follow the same photodestruction law regardless of metallicity. We then compare these results with observations of starburst galaxies, H ii galaxies, and blue compact dwarf galaxies (BCDs). While the integrated mid-infrared spectra of BCDs are reminiscent of a warm dusty ionized gas, we observe a significant contribution to the PAH emission in starburst galaxies that is not arising from PDRs.

Published in ApJ Available from arXiv:1012.1728

Nearby galaxies in more distant contexts

Michael Eskew^{1} and Dennis Zaritsky^{1}

1Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721, USA

We use published reconstructions of the star formation history (SFH) of the Large Magellanic Cloud (LMC), Small Magellanic Cloud, and NGC 300 from the analysis of resolved stellar populations to investigate where such galaxies might land on well-known extragalactic diagnostic plots over the galaxies’ lifetime (assuming that nothing other than their stellar populations change). For example, we find that the evolution of these galaxies implies a complex evolution in the Tully–Fisher relation with lookback time and that the observed scatter is consistent with excursions these galaxies take as their stellar populations evolve. We find that the growth of stellar mass is weighted to early times, despite the strongly star-forming current nature of the three systems. Lastly, we find that these galaxies can take circuitous paths across the color-magnitude diagram. For example, it is possible, within the constraints provided by the current determination of its SFH, that the LMC reached the red sequence at intermediate age prior to ending back up on the blue cloud at the current time. Unfortunately, this behavior happens at sufficiently early times that our resolved SFH is crude and insufficiently constraining to convincingly demonstrate that this was the actual evolutionary path. The limited sample size precludes any general conclusions, but we present these as examples how we can bridge the study of resolved populations and the more distant universe.

Published in The Astronomical Journal Available from arXiv:1101.2262

Conference Papers

Feedback from massive YSOs and massive stars

You-Hua Chu^{1} and Robert A. Gruendl^{1}

1University of Illinois at Urbana-Champaign, USA

Massive stars are powerful sources of radiation, stellar winds, and supernova explosions. The radiative and mechanical energies injected by massive stars into the interstellar medium (ISM) profoundly alter the structure and evolution of the ISM, which subsequently influences the star formation and chemical evolution of the host galaxy. In this review, we will use the Large Magellanic Cloud (LMC) as a laboratory to showcase effects of energy feedback from massive young stellar objects (YSOs) and mature stars. We will also use the Carina Nebula in the Galaxy to illustrate a multi-wavelength study of feedback from massive star.

Oral contribution, published in ”The Multi-wavelength View of Hot, Massive Stars”, Liège, Belgium, July 2010 Available from arXiv:1012.1373

Life and death of very massive stars

Norhasliza Yusof^{1,2}, Hasan Abu Kassim^{1}, Raphael Hirschi^{2,3}, Paul. A. Crowther^{4}, Olivier Schnurr^{5}, Richard J. Parker^{6} and Simon P. Goodwin^{4}

1Department of Physics, University of Malaya, 50603 Kuala Lumpur, Malaysia 2Astrophysics Group, EPSAM, University of Keele, Lennard-Jones Lab, Keele, ST5 5BG, UK 3Institute for the Physics and Mathematics of the Universe, University of Tokyo, 5-1-5 Kashiwanoha, Kashiwa, 277-8583, Japan 4Department of Physics and Astronomy, University of Sheffield, Sheffield S3 7RH, UK 5Astrophysikalisches Institut of Postdam, An der Sternwarte 16, D-14482 Postdam, Germany 6Institute for Astronomy, ETH Zurich, Wolfgang-Pauli-Straße 27, 8093 Zurich, Switzerland

We recently determined the mass of the most massive star known to the date, R 136a1 with a mass at birth 320 times the mass of our sun, as well as the mass of several other stars that are more massive than 150 M_{⊙}. Such massive stars (∼150–300 M_{⊙}) may end their life as pair-instability supernovae (PISN) if they retain enough mass until they die. We have calculated a grid of stellar evolution models in order to investigate the impact of mass loss and rotation on the evolution and fate of these very massive stars. As mass loss is very strong at solar metallicity, our models predict that most of the very massive stars will die as type Ic SNe. Only slowly and non-rotating stars at metallicities below that of the LMC might retain enough mass to produce a PISN. This would mean that the first stellar generations might have produced PISN although their chemical signature is not observed in extremely metal poor stars in the halo of our galaxy.

Poster contribution, published in Science Available from arXiv:1012.3649

The most luminous stars in the Galaxy and the Magellanic Clouds

Wolf-Rainer Hamann^{1}, Andreas Barniske^{1}, Adriane Liermann^{1}, Lidia M. Oskinova^{1}, Diana Pasemann^{1} and Ute Rühling^{1}

1Universität Potsdam, Germany

Some of the Wolf–Rayet (WR) stars are found to have very high bolometric luminosities (log L/L_{⊙} > 6). We employ the Potsdam Wolf–Rayet (PoWR) model atmospheres for their spectral analysis, which yields the bolometric corrections. Distance and interstellar reddening also enter the luminosity estimates. Among the Galactic stars, there is a group of very luminous WNL stars (i.e. WR stars of late subtype from nitrogen sequence with hydrogen being depleted in their atmospheres, but not absent). Their distances are often the major source of uncertainty. From K-band spectroscopy we found a very luminous star (log L/L_{⊙} = 6.5) in the Galactic center region, which we termed the Peony Star because of the form of its surrounding dusty nebula. A similar group of very luminous WNL stars is found in the Large Magellanic Cloud (LMC). In the Small Magellanic Cloud (SMC) the majority of WR stars resides in binary systems. The single WNL stars in the SMC are not very luminous. We conclude that a significant number of very luminous WNL stars exist in the Galaxy and the LMC. With initial masses above 60 M_{⊙}, they apparently evolved directly to the WNL stage without a prior excursion to the red side of the HRD. At the low metallicity of the SMC, the binary channel may be dominant for the formation of WR stars.

Oral contribution, published in ”The multi-wavelength view of hot, massive stars”, Liége 2010 Available from arXiv:1012.1875 and from http://www.astro.physik.uni-potsdam.de/∼www/research/abstracts/wrhamann-liege2010.html

Rubidium-rich Asymptotic Giant Branch stars in the Magellanic Clouds

D.A. García-Hernández^{(1,2)}

1IAC, Spain 2ULL, Spain

The Magellanic Clouds (MCs) offer a unique opportunity to study the stellar evolution and nucleosynthesis of massive Asymptotic Giant Branch (AGB) stars in low metallicity environments where distances are known. Rubidium is a key element to distinguish between high mass AGB stars and low mass AGBs or other type of astronomical objects such as massive red supergiant stars. Theoretically, high mass AGBs are predicted to produce a lot of Rb. We present the discovery of massive Rb-rich AGB stars in the MCs, confirming for the first time that these stars also exist in other galaxies. Our findings show that these stars are generally brighter than the standard adopted luminosity limit (M_{bol} ∼ −7.1) for AGB stars. The observations of massive MC AGBs are qualitatively predicted by the present theoretical models. However, these theoretical models are far from matching the extremely high Rb overabundances observed. This might be related with an incomplete present understanding of the atmospheres of these stars.

Oral contribution, published in ”Why galaxies care about AGB stars II” Available from arXiv:1101.3515

Job Advert

Postdoctoral position in resolved stellar populations

Applications are invited for a postdoctoral position at the University of Tasmania in Hobart, Australia. The research fellow will work with Dr. Andrew Cole on spectroscopic and photometric studies of the Magellanic Clouds to characterize their chemical evolution, kinematics, and star formation histories. The successful applicant will work with project collaborators at the Indian Institute of Astrophysics and will have the opportunity to conduct personal research using the UTAS 1.3m optical telescope that will see first light in 2011, and other astronomical facilities.

The applicant must have a Ph.D. in astronomy, physics, or a closely related field, and a track record of original research in stellar populations, nearby galaxies, and/or multiobject spectroscopy. The appointment is supported by a grant from the Australia–India Strategic Research Fund and will be for a term of two years commencing 1 July or as soon as possible thereafter. The application closing deadline is February 28, 2011. Enquiries may be addressed to andrew.cole@utas.edu.au; more information is also available from the URL listed below. Applications must be made via the webpage listed below.

See also http://jobs.admin.utas.edu.au/positions/pd detail.aspx?PositionId=1723

Announcement

MAGMA LMC Early Data Release

The Magellanic Mopra Assessment (MAGMA), which has undertaken a CO mapping survey of the LMC using the Mopra Telescope, has made available an early data release (EDR) image on January 11, to coincide with the winter AAS meeting in Seattle. The image (a CO integrated intensity map) is available on the MAGMA website at http://mmwave.astro.illinois.edu/magma/. Please read the accompanying text file for information about how the image was produced. Users may cite T. Wong et al., 2011, in preparation, as well as A. Hughes et al., MNRAS 406, 2065 (2010).

See also http://mmwave.astro.illinois.edu/magma/